OpenStudy (anonymous):

if y varies directly as x and y = 540 when x= 10, find x when y = 1080

OpenStudy (anonymous):

y = kx

OpenStudy (anonymous):

y = kx is the formula for direct variation

OpenStudy (anonymous):

540 = 10k

OpenStudy (anonymous):

54 = k

OpenStudy (anonymous):

1080 = 54x

OpenStudy (anonymous):

1080/54 = x

OpenStudy (anonymous):

x =20

OpenStudy (anonymous):

540 = k(10) Divide both sides by 10 54 = k

OpenStudy (anonymous):

1080 = 54 x 1080/54 = x

OpenStudy (anonymous):

x = 20

OpenStudy (anonymous):

The equation for Direct Variation is y=kx